OpenStudy (anonymous):

If the radius and height of a cylinder are multiplied by 4, the lateral area of the cylinder is multiplied by 2. 4. 8. 16.

zepdrix (zepdrix):

The `Lateral Area of a Cylinder` is given by, \(\large A=2\pi r h\) They're asking us: If we replace \(\large \color{orangered}{r}\) with \(\large \color{orangered}{4r}\), and \(\large \color{green}{h}\) with \(\large \color{green}{4h}\), What happens to \(\large A\)? \(\large A=2\pi\color{orangered}{r}\color{green}{h} \qquad \rightarrow \qquad A=2\pi\color{orangered}{(4r)}\color{green}{(4h)}\) What is the difference between our first A and our second one? :)